Circular No. 7153

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SAGITTARII 1999 Y. and R. Kushida, Yatsugatake South Base Observatory, report the

discovery by Minoru Yamamoto, Okazaki, Aichi, Japan, of a nova (mag 8.6) on two TMax-400 films taken with a 200-mm f/4 camera lens (+ PO-0 filter) on Apr. 25.731 UT. No star is visible at this location on 23 previous films taken by Yamamoto dating back to 1997 Feb. 19, the most recent from 1999 Apr. 11 (limiting mag 10.6). The Kushidas provide the following precise position from an unfiltered CCD image taken with a 0.4-m f/6.3 Schmidt-Cassegrain reflector on Apr. 26.778, when the star appeared at mag 7.6: α = 18^{h}07^{m}36^{s}.22, δ = −27^{o}20′13′′.5 (equinox 2000.0). Nothing obvious appears at this location on the Digital Sky Survey.

W. Liller, Viña del Mar, Chile, writes: “A low-resolution CCD spectrogram of Yamamoto’s new star obtained on Apr. 27.195 UT with a 0.20-m Schmidt camera (+ objective prism) shows Hα weakly in emission, ∼ 20 percent brighter than the neighboring continuum, indicating that it is a nova in the early stages of development. On Apr. 27.18, its CCD magnitudes was V = 8.45 (B9 comparison star SAO 186331, Tycho V = 8.58).”

COMET C/1999 H1 (LEE) Ephemeris extension to IAUC 7147 (elements on MPEC 1999-H33):

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Apr. 22 12^{h}17.^{m}66 −65^{o}52.^{′}7 0.844 1.630 123.^{o}5 30.^{o}9 8.8
24 11 45.55 −63 49.5 0.811 1.602 123.4 31.6 8.6
26 11 16.54 −61 09.0 0.782 1.573 122.7 32.6 8.4
28 10 51.15 −57 54.4 0.759 1.544 121.4 33.8 8.3
30 10 29.40 −54 10.6 0.741 1.515 119.4 35.4 8.2
May 2 10 10.98 −50 03.6 0.728 1.486 116.9 37.2 8.0
4 9 55.46 −45 40.0 0.721 1.458 113.9 39.2 7.9
6 9 42.39 −41 06.7 0.720 1.429 110.4 41.4 7.8
8 9 31.36 −36 30.3 0.725 1.400 106.5 43.7 7.8
10 9 22.02 −31 56.8 0.735 1.371 102.5 46.0 7.7
12 9 14.07 −27 31.1 0.750 1.342  98.3 48.1 7.7
14 9 07.27 −23 16.9 0.770 1.313  94.1 50.1 7.6
16 9 01.42 −19 16.8 0.794 1.284  89.9 51.9 7.6
18 8 56.35 −15 32.0 0.822 1.256  85.8 53.4 7.6
20 8 51.94 −12 03.1 0.852 1.227  81.9 54.7 7.5
22 8 48.07 − 8 49.7 0.885 1.199  78.0 55.7 7.5
1999 April 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green